The velocity potential function ϕ is given by ϕ = x2 - y^2. Find the velocity components in x and y direction. Also, show that ϕ represents a possible case of fluid flow.


1
Expert's answer
2022-02-15T09:58:22-0500

In velocity we can write 



"u=\dfrac{\partial\varphi}{\partial x}, v=\frac{\partial\varphi}{\partial y}"


Given "\varphi=x^2-y^2"

"u=\dfrac{\partial\varphi}{\partial x}=2x, v=\dfrac{\partial\varphi}{\partial y}=-2y"




"\dfrac{\partial u}{\partial x}=\dfrac{\partial^2\varphi}{\partial x^2}=2""\dfrac{\partial v}{\partial y}=\dfrac{\partial^2\varphi}{\partial y^2}==2""\dfrac{\partial u}{\partial x}+\dfrac{\partial v}{\partial y}=0"

The continuity equation is satisfied and the flow is irrotational.